# Implementing LLMs in Industrial Process modeling: Addressing categorical variables

### v1

@@ 5.4 Mean squared error, MSE
Then, in the first case (see Fig. 8), we explore the original model, which shows an increasing trend as the training data size grows (dashed black line). However, its confidence interval also widens downward once 90% of the training data is considered, which means it limits the potential for improvement in this metric, as the bounds of the other (blue and green) confidence intervals within which the true value of is likely to fall are slightly above those of the original model.
In contrast, when examining the curves for models using Doc2Vec to encode only the shapes of the inserts (in blue) or both the shapes of the inserts and the recipe (in green), the average follows a similar increasing trend, slightly above the black curve. Additionally, their respective confidence intervals also widen once 90% of the data is included.
In the remaining two cases (see Fig. 8), similar behaviors are observed, although the average scores in each case are closer to the original model’s average , indicating no significant improvements. For transformers, the same widening of confidence intervals is noted when considering 90% of the training data, suggesting that the model requires as much data as possible to achieve only slight improvements in performance. Again, this approach helps to reduce the variability of the MSE mean, ensuring that the true value falls within this interval.
@@
Finally, to analyze the performance of the XGBOOST model when varying the dimensionality of the latent spaces generated by transformers and embedding, we also estimated the confidence intervals for the average MSE and . Considering much larger latent space dimensions for the Doc2Vec model do not show significant improvements compared to those already achieved with a dimension of 3. However, it is more interesting to observe the effects when reducing the latent space dimension of the transformers to 3 (cf. Appendix A, Fig. 15 and Fig. 16). The main advantages here are a slight improvement in metrics and increased efficiency of the predictive model, as it has to handle fewer variables.
@@ 5.6 Feature importance and Shapley Analysis
For the original model (see Fig. 9), which uses binary encoding for handling categorical features, we observe that numerical features (in blue) are the most influential in terms of feature importance scores, particularly those related to total area and surface variability. In contrast, categorical variables related to insert geometries (in yellow) and the recipe (in green) have considerably lower scores. Shapley analysis shows a similar trend where numerical features have higher scores than categorical ones. This suggests that potentially critical inputs for the process are neglected as categorical variables.
Both feature importance and Shapley analysis suggest that, although some numerical features remain the most important, we observe a slight increase in the importance of insert shapes after being encoded with Doc2Vec (see Fig. 10). It is also important to note that the number of input features has tripled following preprocessing and encoding. This indicates that incorporating more useful information into the regression model is indeed possible by transforming the descriptions of insert shapes using an embedding.
Finally, when we use more sophisticated models as transformers to encode categorical features (see Figs. 11 and 12), a similar trend continues in terms of feature importance and Shapley analysis: categorical variables related to the shapes of the inserts are ranked within the top 5. This further emphasizes the value of having interpreted textual descriptions more robustly for inclusion in the regression model.
@@ 6 Conclusions
This case study highlights the significant benefits derived from leveraging meaningful embeddings, particularly through the use of transformer models, for processing categorical variables. Transformers, known for their ability to capture intricate patterns and relationships in data, enable categorical variables to be represented in a manner that captures their semantic and contextual meaning. By doing so, the embeddings generated reflect not only the distinct values of these categorical variables but also the underlying relationships between them, which might otherwise be missed by traditional encoding techniques such as one-hot encoding.
The ability of transformers to handle categorical variables more effectively results in a substantial improvement in the performance and efficiency of predictive models. By embedding categorical data in a way that is contextually rich and semantically meaningful, the model can extract deeper insights and make more accurate predictions. This approach ensures that models are better equipped to handle complex, real-world datasets where categorical variables often play a crucial role in determining outcomes.
In addition to enhancing predictive capabilities, this method improves the assessment of critical process parameters. Categorical variables can now be considered in a manner that is more representative of actual observations and more aligned with the way data behaves in real-world scenarios. This leads to a more nuanced understanding of the interactions between variables and offers a more comprehensive approach to process optimization. By embedding categorical data in a meaningful way, models can incorporate these variables into their structure more naturally, enhancing both the interpretability and the accuracy of the results.
Looking forward, there is great potential in expanding this work into sensitivity analysis and uncertainty quantification. These are critical areas for model evaluation, as they provide insights into how changes in input variables affect the output and help quantify the uncertainty associated with predictions. Future research could focus on exploring more advanced machine learning methods and alternative mathematical frameworks to further improve both the accuracy and interpretability of models. Methods such as Bayesian neural networks, variational inference, and probabilistic graphical models could provide robust frameworks for uncertainty quantification, offering deeper insights into the reliability of predictions.
Moreover, by integrating these advanced techniques into the existing framework for processing categorical variables, future studies could refine sensitivity analysis, making it more adaptable to complex, high-dimensional data. This would not only boost the performance of models but also provide more granular control over how categorical data influences predictions, enabling more informed decision-making in critical applications.
@@ Appendix A Appendix
A.1 Similarity after dimensionality Reduction
A.2 XGBOOST performance after dimensionality Reduction

### v2

@@ 5.5 R2 score
Thus, in Fig. 9, we explore the original model performance guides by the score, which shows an increasing trend as the training data size grows (dashed black line). Its confidence interval also widens downward once 90% of the training data is considered, which means it limits the potential for improvement in this metric, as the bounds of the other (blue and green) confidence intervals within which the true value of is likely to fall are slightly above those of the original model.
In contrast, when examining the curves for models using Doc2Vec to encode only the Insert shape features (in blue) or both the shapes of the inserts and the recipe (in green), the average follows a similar increasing trend, slightly above the black curve. Additionally, their respective confidence intervals also widen once 90% of the data is included.
In the remaining two cases, corresponding to the implementation of all-MiniLM-L12-v2 (Fig. 9 (b) and (b’)) and all-mpnet-base-v2 (Fig. 9 (c) and (c’)) transformers with and without the implementation of PCA, respectively), similar behaviors are observed, although the average scores in each case are closer to the original model’s average , indicating no significant improvements. For transformers, the same widening of confidence intervals is noted when considering 90% of the training data, suggesting that the model requires as much data as possible to achieve only slight improvements in performance. Again, this approach helps to reduce the variability of the mean, ensuring that the true value falls within this interval.
@@
Finally, to analyze the performance of the XGBOOST model when varying the dimensionality of the latent spaces generated by transformers and embedding, we also estimated the confidence intervals for the average MSE and . Considering much larger latent space dimensions for the Doc2Vec model do not show significant improvements compared to those already achieved with a dimension of 3. However, it is more interesting to observe the effects when reducing the latent space dimension of the transformers to 3 (cf. Figs. 8 (b’) and (c’), 9 (b’) and (c’), 15, and 16). The main advantages here are a slight improvement in metrics and increased efficiency of the predictive model, as it has to handle fewer variables.
Another important aspect to mention at this point is the computational cost of each of the models presented. First, it is worth noting that there is no significant computational cost associated with training the Doc2Vec embedding due to the small quantity and short length of the descriptions, as well as the fact that the latent space dimension, set to 3, is low. Additionally, the computational cost of applying the two transformers (all-MiniLM-L12-v2 and all-MPNet-base-v2) is negligible because both are pre-trained models and their respective encoding speeds are highly efficient, as reported in Table 6 (cf. [40]).
| Model | # Parameters | Encoding speed |
| (sentences/sec) | ||
| Doc2Vec | for this application | 50K100K |
| all-MiniLM-L12-v2 | 33.4M | 7.5K |
| all-mpnet-base-v2 | 109M | 2.8K |
The computational cost arises when training the predictive model, which incorporates the vector representations of the insert shapes descriptions using embedding and transformers. Each model is trained six times, once for each dataset proportion (see Fig. 8 and Fig. 9). Thus, each of the predictive models using the vector representation obtained through the Doc2Vec embedding takes, on average, 23 seconds per run. In contrast, the predictive models that include high-dimensional vector representations produced by the transformers have a much higher computational cost. The all-MiniLM-L12-v2 model takes, on average, about 10 minutes per run, while the all-mpnet-base-v2 model can take approximately 20 minutes per run. This demonstrates a clear correlation between dimensionality and the computational cost of each model.
However, this problem caused by high dimensionality can be successfully addressed by reducing the dimensionality of the vector representations with methods such as PCA or UMAP, as detailed in Section 5.2 and shown in Section 5.4, Section 5.5 and Appendix A. This approach yields results with high precision and an average runtime of around 25 seconds, thus recovering efficiency without losing accuracy.
@@ 5.6 Feature importance and Shapley Analysis
Additionally, it is important to mention that these values do not follow to a standardized range, and there is no upper limit to the total_gain score or the Shapley scores, as we are summing the overall contributions of the feature. As a result, the total_gain score can potentially increase indefinitely, reflecting the cumulative impact of all contributing factors.
For the original model (see Fig. 10), which uses binary encoding for handling categorical features, we observe that numerical features (in gray) are the most influential in terms of feature importance scores, particularly those related to total area and surface variability. In contrast, categorical variables related to insert geometries (in blue) and the recipe (in black) have considerably lower scores. Shapley analysis shows a similar trend where numerical features have higher scores than categorical ones. This suggests that potentially critical inputs for the process are neglected as categorical variables.
Based on subject matter expertise, we recognize the significance of categorical variables; however, quantifying this importance directly in the original model is not possible. With the implementation of embeddings, we can now effectively quantify feature importance. Furthermore, we observe that categorical features, after being transformed through embeddings, have begun to dominate and become increasingly significant.
Both feature importance and Shapley analysis suggest that, although some numerical features remain the most important, we observe a slight increase in the importance of insert shapes after being encoded with Doc2Vec (see Fig. 11). It is also important to note that the number of input features has tripled following preprocessing and encoding. This indicates that incorporating more useful information into the predictive model is indeed possible by transforming the descriptions of insert shapes using an embedding.
The use of more sophisticated models as transformers to encode categorical features (see Fig. 12 and Fig. 13), show a similar trend in terms of feature importance and Shapley analysis: categorical variables related to the shapes of the inserts are ranked within the top 5. This further emphasizes the value of having interpreted textual descriptions more robustly for inclusion in the predictive model.
Finally, with this set of tools, we can initiate the optimization of the reactor setup. An initial step involves leveraging the above results to identify the most influential input parameters for the predictive model. This analysis can be further refined through sensitivity analysis and uncertainty quantification, enabling a deeper understanding of how these parameters contribute to process variability. By systematically addressing these factors, we aim to mitigate variability and enhance process stability. This effort remains an integral part of our ongoing research.
@@ 6 Conclusions
This study highlights the benefits of leveraging transformer-based embeddings for processing categorical variables, particularly in industrial applications. The case study presented here, concerns the industrial-scale Chemical Vapour Deposition of coatings, involving actual production data. The nature of the inputs that include both numerical and categorical industrial production data, makes this a suitable example to showcase the performance of the natural language processing models. Nevertheless the findings of this work are not limited to this specific application but can be useful in other cases where there is variety in the type of data involved. Unlike traditional encoding techniques such as one-hot or binary encoding, transformers capture intricate patterns and relationships within categorical data, enabling representations that preserve both semantic meaning and contextual relevance.
By embedding categorical data in a contextually rich manner, models gain deeper insights, leading to a more comprehensive understanding of critical process parameters. This approach is especially valuable in complex industrial systems, where categorical inputs play a pivotal role in determining outcomes.
While this study focuses on the methodological advantages of transformer-based embeddings, future research will explore their direct applications in reactor optimization, process efficiency improvements, and the development of novel operational strategies. These extensions present distinct research challenges and merit a separate investigation to ensure a thorough and application-driven analysis.
Furthermore, this work lays the foundation for incorporating sensitivity analysis and uncertainty quantification into model evaluation. Future studies could explore probabilistic frameworks such as Bayesian neural networks, variational inference, and probabilistic graphical models to enhance both accuracy and interpretability. By integrating these techniques with transformer-based embeddings, future research can refine sensitivity analysis for complex, high-dimensional data, ultimately improving decision-making in industrial applications.
By demonstrating the advantages of embedding categorical variables using transformers, this work contributes to a broader understanding of how NLP-based techniques can be applied in industrial machine learning. The insights gained serve as a foundation for further exploration, paving the way for more robust, interpretable, and application-driven advancements in process modeling and optimization.
@@ Appendix A Appendix
A.1 Similarity after dimensionality reduction using UMAP
A.2 XGBOOST performance after dimensionality reduction using UMAP